OpenStudy (anonymous):

Can anyone explain how to do this problem: A grinding wheel .29m in diameter spins up from rest to 2705 rpm in 2 seconds a- Calculate its final angular velocity in rad/s b- what is the angular acceleration of the grinding wheel c- what is the linear speed of a point on the edge of the grinding wheel after 2 seconds d- what is the total linear acceleration of a point on the edge of the grinding wheel after 2 seconds e- if the grinding wheel has a mass of 3kg and can be considered as a uniform soild disk what is the magnitude of the torque acting on the grinding wheel's angular mome

OpenStudy (anonymous):

\[\frac{ rev.s }{ \min} *\frac{ 1 \min }{ 60 \sec } * \frac{ 2\pi rad.s }{ rev} =\frac{ rad.s }{ \sec }\]

OpenStudy (anonymous):

can you do 'a' now?

OpenStudy (anonymous):

YES! thank you, I got A now ....can you help me solve any of the others?

OpenStudy (anonymous):

sure... so you have omega f and omega f = omega i + alpha *t

OpenStudy (anonymous):

linear speed is V = r*omega

OpenStudy (anonymous):

good on those two?

OpenStudy (anonymous):

still a bit puzzled on "B"

OpenStudy (anonymous):

so I'm still figuring out how to write it with the numbers given to me

OpenStudy (anonymous):

\[\omega _{f} = \omega _{i} + \alpha t\]

OpenStudy (anonymous):

it's just like the equation from kinematics... remember? Vf = Vi + at

OpenStudy (anonymous):

Final velocity = to initial velocity + acceleration * Time?

OpenStudy (anonymous):

here the quantities are all angular(except time of course) ... angular velocity, angular acceleration etc.

OpenStudy (anonymous):

yep, that one.

OpenStudy (anonymous):

okay now its coming to me a bit so final w is = to initial w + acceleration * time? "w" would stand for angular?

OpenStudy (anonymous):

yes \[\omega \] is angular velocity \[\alpha \] is angular acceleration \[\theta \] is angular displacement

OpenStudy (anonymous):

so here I'm gonna draw out my equation , which hopefully is write

OpenStudy (anonymous):

right*

OpenStudy (anonymous):

I got it now

OpenStudy (anonymous):

so now I'm confused about c

OpenStudy (anonymous):

what about it?

OpenStudy (anonymous):

I know that the speed of the edge of the grinding wheel will be faster ,but really can't remember how

OpenStudy (anonymous):

I just can't figure out the formula to use

OpenStudy (anonymous):

V=r*omega

OpenStudy (anonymous):

but in this case the number are .29m, 2705 rpm and 2 seconds

OpenStudy (anonymous):

.29m is diameter, not radius. we don't use rpm's we use rad.s/sec.

OpenStudy (anonymous):

okay I think I should get this, might take me a while I'll message back if I need any help thanks

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

if the trampoline behaves like a spring with a spring stiffness constant k= 3 X 10^4 n/m , how far does the person depress it? I know the height = 3 and his speed as he lands on the trampoline is 10.38 m/s. What is the equation?